zarkam21:

Equilibrium? (Help with concept) Finding eqb visually/graphically , Finding eqb algebraically??

Vocaloid:

what exactly do you mean by "equilibrium"?

zarkam21:

Idk just a concept my professor said to go over and practice graphically algebraically

zarkam21:

I'm in Calc I if that helps

zarkam21:

LIke I said I'm having my midterm tomorrow so just going over stuff and the list said equilibrium which I do not understand. So the best you can explain will help

Vocaloid:

i've not really heard this term before in the context of calculus do you have an example from your lecture notes or textbook?

zarkam21:

1 attachment
zarkam21:

equilibria?

Vocaloid:

oh ok I got it, you just need to solve when x_n+1 = x_n (they can use the variable x, or a, or whatever)

Vocaloid:

like, for the first example a_n+1 = sqrt(a) you just set a = a_n+1 to get a = sqrt(a) which is only true for a = 1 or 0

Vocaloid:

it's just algebra, really

zarkam21:

the n+1 disappears?

Vocaloid:

we substitute a for a_n+1 since a = a_n+1

Vocaloid:

if it makes it less confusing you can just use a new variable like z or something a_n+1 = sqrt(a) becomes z = sqrt(z) which gives the same result, just make sure to report your answer in terms of a_n or whatever variable is already in the problem

zarkam21:

okay so what if numbers were involved?

zarkam21:

same concept

zarkam21:

?

Vocaloid:

yeah, you just set a_n+1 and a_n equal and solve, regardless of what else is in the equation

zarkam21:

can you just give a quick example with numbers ? :S

zarkam21:

Just want to note an example down for reference with actuAl numbers

Vocaloid:

like, the second example you have has a_n+1 = 0.33an + 214.5 you just set a_n+1 and an equal to each other to get a = 0.33a + 214.5 and solve for a

zarkam21:

a-214.5=0.33a

Vocaloid:

I would recommend subtracting 0.33a from both sides and dividing to isolate a

zarkam21:

a-0.33a=214.5

Vocaloid:

good, finish simplifying the left side

zarkam21:

6.71E-1a=214.5 a=3.2E2

Vocaloid:

I would recommend not converting to base 10 notation here

zarkam21:

.67 do 214.5/.67=320

Vocaloid:

like, a-0.33a is just 1 - 0.33 which is 0.67a 0.67a = 214.5 a = 214.5/0.67 = 320.149

Vocaloid:

yes

zarkam21:

NOw how about to graphing an equilibria or whatever :S

Vocaloid:

uh well since a_n+1 = an you should look for a part of the graph that levels off/flatlines

Vocaloid:

i think